### v1

# Systematic Evaluation of LLM-as-a-Judge in LLM Alignment Tasks: Explainable Metrics and Diverse Prompt Templates

Abstract
Alignment approaches such as RLHF and DPO are actively investigated to align large language models (LLMs) with human preferences. Commercial large language models (LLMs) like GPT-4 have been recently employed to evaluate and compare different LLM alignment approaches. These models act as surrogates for human evaluators due to their promising abilities to approximate human preferences with remarkably faster feedback and lower costs. This methodology is referred to as LLM-as-a-judge. However, concerns regarding its reliability have emerged, attributed to LLM judges’ biases and inconsistent decision-making. Previous research has sought to develop robust evaluation frameworks for assessing the reliability of LLM judges and their alignment with human preferences. However, the employed evaluation metrics often lack adequate explainability and fail to address the internal inconsistency of LLMs. Additionally, existing studies inadequately explore the impact of various prompt templates when applying LLM-as-a-judge methods, which leads to potentially inconsistent comparisons between different alignment algorithms. In this work, we systematically evaluate LLM judges on alignment tasks (e.g. summarization) by defining evaluation metrics with improved theoretical interpretability and disentangling reliability metrics with LLM internal inconsistency. We develop a framework to evaluate, compare, and visualize the reliability and alignment of LLM judges to provide informative observations that help choose LLM judges for alignment tasks. In the experiments, we examine the effect of diverse prompt templates on LLM-judge reliability and also demonstrate our developed framework by evaluating and comparing various LLM judges on two common alignment datasets. Our results indicate a significant impact of prompt templates on LLM judge performance, as well as a mediocre alignment level between the tested LLM judges and human evaluators. The code of the developed framework is available at https://github.com/shenghh2015/llm-judge-eval.
**footnotetext: Equal contribution; corresponding authors:footnotetext: 🖂 huiwei@cs.umass.edu,  🖂 shenghh2015@gmail.comfootnotetext: Work was done when H. Wei & J. Lin were interns at PAII Inc.
Introduction
Alignment techniques, such as RLHF [1, 2, 3, 4], DPO [5, 6], and N-best sampling (or rejection sampling) [7, 8] have been actively investigated to align large language models (LLMs) with human preferences. These studies typically use human-based pairwise evaluations as the gold standard for method evaluation and comparison. During the evaluation procedure, the human judge is presented with a question and two associated responses generated by different LLMs and is tasked with evaluating which response is preferred based on general criteria, such as helpfulness, honesty, and harmlessness [9]. A win-rate metric is subsequently calculated based on these judgment results and utilized to assess which LLM more effectively aligns with human preferences. Despite its high effectiveness, human-based evaluation is notably slow and costly [10], rendering it generally impractical for rapid assessments and advancements in alignment methodologies.
Recently, commercial LLMs, such as GPT-4 [11] and GPT-3.5-turbo [12], have been widely used as the surrogates for human evaluators, referred to as LLM-as-a-judge, to perform pairwise evaluation on numerous LLM alignment tasks, such as summarization [5, 13, 14, 15] as well as single- or multi-turn conversations [5, 13, 15, 16, 17, 18]. Since these commercial models have already been extensively trained with advanced alignment techniques [11, 4], they are promisingly capable of approximating human preferences [5, 19].
While it is plausible to utilize these models as surrogates for human judges, biases and inconsistencies are frequently observed in their judgment results, despite the application of various bias-mitigation techniques[20, 5]. This necessitates a systematic investigation of LLM judge reliability and alignment with human preferences in the context of LLM alignment tasks.
Previous studies have evaluated LLM-as-a-judge methods on various language generation tasks [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 19, 33, 34, 35]. However, these studies encounter three main limitations:
  1. 1.
    Lacking clear theoretical interpretability for bias definitions (e.g. position bias and length bias).
  2. 2.
    Not considering internal inconsistencies (i.e., system noise) by assuming LLM judges make deterministic decisions across identical experiments.
  3. 3.
    Concentrating on evaluating various LLMs, while the effects of prompt templates have been insufficiently examined.
In this study, we aim to address these limitations and advance the systematic evaluation of LLM judges on LLM alignment tasks. The contributions of our study are:
  1. 1.
    We enhance the theoretical explainability of current evaluation metrics to assess LLM-judge alignment with human preferences and their reliability, including accuracy, position bias, and length bias by defining them within a unified evaluation framework. Then, we provide practical ways to compute these metrics. In addition, we explicitly define and measure the LLM internal self-inconsistency as flipping noise, and mitigate its impact on position bias and length bias. To the best of our knowledge, this is the first study to address this issue.
  2. 2.
    We develop a framework to evaluate, compare, and visualize the alignment and reliability of LLM judges, with a general and flexible design, allowing for application across a wide range of LLMs and user-defined prompt templates. We utilize a wide range of up-to-date prompt templates with diverse formats to investigate their impact on LLM judge performance. We also demonstrate our proposed framework through experiments to evaluate and compare various LLM judges comprehensively and consistently.
  3. 3.
    Our results indicate a significant impact of prompt templates on LLM judge performance, as well as a mediocre alignment level between the tested LLM judges and human evaluators. This underscores the need for a thorough and careful comparison of various LLMs and prompt templates before employing the LLM-as-a-judge methodology. Additionally, it highlights the importance of human evaluation for achieving more precise comparisons between different LLM alignment systems, provided that time and cost constraints are manageable. Finally, we present the ranking results of all the LLM judges for both datasets to facilitate the selection of the most appropriate judge.
Background and Related Work
In this section, we define the pairwise evaluation task conducted by both human and LLM judges, and examine self-inconsistencies and biases inherent in LLM judges. Additionally, we review relevant literature on position bias and length bias.
Human-based Pairwise Evaluation
Given a set of $N$ questions, each paired with responses generated by separate LLMs, the human judge is asked to select the better response based on predefined criteria, such as coherence and helpfulness, as specified in the introduction. Let $N_{1}$ and $N_{2}$ be the numbers that the first and second answer are chosen. The win rate of the first and the second LLM is defined as $w_{1,2}\!=\!\frac{N_{1}}{N}$ and $w_{2,1}\!=\!1\!-\!w_{1,2}$, respectively.
LLM-based Pairwise Evaluation
LLM-judges are subjected to the same evaluation procedures as human judges. However, compared with humans, LLMs are more sensitive to instructions (i.e. prompt templates) [36, 30]. Thus, in this study, we define an LLM-judge as the combination of a specific LLM and a particular prompt template.
LLM-Judge Self-Inconsistency
Previous studies have observed that LLM judges [25, 36] may produce inconsistent judgments even when presented with identical prompts. This is caused by non-greedy decoding strategies leveraged by LLMs, such as top-p and top-k, which generate non-deterministic outputs. The non-deterministic level is controlled by the parameter temperature. In this work, we refer to these inconsistencies as self-inconsistency or system noise in LLM judges and model and quantify them using the concept of flipping noise.
LLM-Judge Bias
Position bias and length bias are two predominant biases frequently observed in LLM judges utilizing commercial LLMs.
Positioin bias refers to LLM-judge’s systematic preference for a specific response position (the first or the second in the pairwise evaluation task). Wang et. al [21] and Lee et. al [37] observed the position bias when using GPT-4 [11] and PaLM 2 [38] as the judge for the pairwise comparison between candidate LLMs. They measured the position bias by the ratio of inconsistent decisions made by LLM judges after swapping response positions. Differently, studies from Liusie et. al [39] and Zheng et. al  [19] quantified the position bias as the disparity of selection probabilities after reversing the response order.
Length bias refers to LLM judge’s systematic preference for longer responses even when their qualities are similar to shorter versions. Saito et al. [22] observed a discrepancy between LLMs and human preferences regarding response length. They employed accuracy parity—related to human preferences for longer responses and shorter responses—to measure relative length bias.
In contrast to the aforementioned studies, our work examines the impact of LLM judge self-inconsistency on the evaluation of both position bias and length bias, and provides methodologies for disentangling these biases from flipping noise. We also offer a theoretical analysis and validation of our defined metrics to enhance their interpretability. Additionally, we investigate the relationship between these biases and accuracy, revealing significant insights. Finally, our study includes an extensive evaluation of position and length bias across a diverse set of LLM judges with various prompt templates.
Methods
In this section, we introduce our evaluation metrics and framework for assessing LLM-judge biases and alignment with human preferences. We begin by introducing notations. Then we present our proposed evaluation metrics with enhanced theoretical interpretability of accuracy, flipping noise, position bias, and length bias, as well as provide practical methods to compute them. Subsequently, we describe our developed framework for the systematic evaluation of LLM-as-a-judge methods.
Notations
Let $\mathcal{D}\!=\!\{h_{n}|n\!=\!1\!\ldots\!N\}$ be a human-preference dataset containing $N$ data cases. An individual data case $h_{n}\!=\!(x^{(n)},y^{(n)}_{c},y^{(n)}_{r})$ represents a prompt-response pair with a human preference label. Here, $x$ is a prompt (e.g. a post for summarization or a question for a question-answering task), $y_{c}$ is the preferred LLM response, and $y_{r}$ is the less preferred response by human evaluators. We assume each case is drawn from the distribution $h_{n}\!\sim\!p(h|\theta)$, where $\theta$ represents the underlying human preferences. We drop the data case index $n$ for brevity when the context is clear.
Evaluation Metrics
Accuracy Accuracy measures the alignment level of LLM judges with human preferences. Formally, we denote $\theta_{l}$ as the underlying preference by some LLM-judge $l$, and accuracy evaluates how closely $\theta_{l}$ is to $\theta$, where $\theta$ is the human preference defined in the last section.
In the literature  [19, 21], there are two versions of the accuracy metric: $\text{Acc}_{\text{both}}$ and $\text{Acc}_{\text{random}}$, and we follow the same definitions in this work. Before formally defining them, we assume the LLM judge decides on each data case by considering two response orders: $h\!=\!(x,y_{c},y_{r})$ and $h^{\prime}\!=\!(x,y_{r},y_{c})$. The judge then selects the preferred response $y$ and $y^{\prime}$ from each order $h$ and $h^{\prime}$, where $y,y^{\prime}\!\in\!\{y_{c},y_{r}\}$. Broadly, we denote the set of judging results as $J\!=\!\{s_{n}|n=1\ldots N\}$, where each result $s_{n}\!=\!(y^{(n)},y^{\prime(n)})$ represents the selection outcome from both response orders across all the data cases in the dataset $\mathcal{D}$. Then the accuracy metrics $\text{Acc}_{\text{both}}$ and $\text{Acc}_{\text{random}}$ can be defined over the judging set $J$ as follows:
$\text{Acc}_{\text{both}}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}\left(y^{(n)}=y^{(n)}_{c}\land y^{\prime(n)}=y^{(n)}_{c}\right),$
$\text{Acc}_{\text{random}}$ $=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}\left(y^{(n)}_{\text{random}}=y^{(n)}_{c}\right),$
where $y_{\text{random}}$ is randomly chosen from $\{y,y^{\prime}\}$ with the probability of 0.5.
Flipping Noise As mentioned in the background section, LLM outputs are generally non-deterministic, which can lead to inconsistent judgments even when the LLM judge is presented with the identical data case $h=(x,y_{c},y_{r})$.
In this section, we model the LLM judge’s decision as a binary variable that indicates if the human-preferred response $y_{c}$ is also selected by the LLM judge. When an inconsistent decision occurs for the same data case, we refer to it as “flipping” the decision to the opposite value. This behavior is quantified using the concept of flipping noise.
Formally, for a data case $(x,y_{c},y_{r})$, we let $X\!\in\!\{0,1\}$ represent if an LLM judge’s choice is aligned with human preferences. Specifically, $X\!=\!1$ denotes that the LLM judge has selected the human preferred response $y_{c}$, whereas $X\!=\!0$ indicates that the alternative response $y_{r}$ has been chosen by the LLM judge. Then we define the LLM judge’s decision after adding the flipping noise (i.e. noisy observation of the decision) using a binary random variable $Z\!\in\!\{0,1\}$ as follows:
$Z$ =$=\begin{cases}1-X,&p\left[1-X|X\right]=q\\ X,&1-q\end{cases}$ (1)
where $q$ is the probability that the LLM judge’s decision is flipped. For a completely deterministic LLM judge, $q\!=\!0$.
Position Bias (PB) As a reminder, we define accuracy based on two sets of responses with reversed orders, namely $(y_{c},y_{r})$ and $(y_{r},y_{c})$, for the same prompt $x$. To assess accuracy, we require the LLM judge to be evaluated in both orders. Here, we employ the same setting to define position bias.
First of all, we define $p\left[X=1|(y_{c},y_{r})\right]$ as the probability that the LLM-judge’s result aligns with the human selection for the response order $(y_{c},y_{r})$, and $p\left[X=1|(y_{c},y_{r})\right]$ as the probability that the LLM-judge’s result aligns with the human selection when the order is reversed. It is important to note these two probabilities are essentially accuracy metrics for the two response positions.
We first consider a special case where the LLM judge makes a fully consistent decision (i.e. $q=0$), and is completely insensitive to the response position order (i.e. exhibits no position bias). This implies that accuracy should be invariant regarding response positions: $p\left[X=1|(y_{c},y_{r})\right]-p\left[X=1|(y_{r},y_{c})\right]=0$.
Additionally, if the LLM-judge exhibits position bias favoring the first position over the second, it will select $y_{c}$ more frequently in $(y_{c},y_{r})$ and $y_{r}$ more frequently in $(y_{c},y_{r})$, compared to the scenario with no position bias. Thus, the accuracy $p\left[X=1|(y_{c},y_{r})\right]$ will increase and $p\left[X\!=\!1|(y_{r},y_{c})\right]$ will decrease, resulting in $p\left[X\!=\!1|(y_{c},y_{r})\right]-p\left[X\!=\!1|(y_{r},y_{c})\right]>0$. The same rationale applies when the second position is preferred.
Based on these intuitions, we define position bias as:
$\text{PB}=p\left[X=1|(y_{c},y_{r})\right]-p\left[X=1|(y_{r},y_{c})\right]$ (2)
[…]
Finally, we address the general case in which the LLM-judge makes non-deterministic decisions and exhibits position bias. Here, only noisy observation $Z$ defined in Eq. 1 is observable, instead of $X$. Thus, to determine the underlying position bias as defined by Eq. 2, we first compute the accuracy of both positions based on $Z$, and then apply the de-noise process according to the following relationships between accuracy based on $X$ and accuracy based on $Z$.
$p\left[X=1|(y_{c},y_{r})\right]$ $=\frac{p\left[Z=1|(y_{c},y_{r})\right]-q_{cr}}{1-2\cdot q_{cr}}$
$p\left[X=1|(y_{r},y_{c})\right]$ $=\frac{p\left[Z=1|(y_{r},y_{c})\right]-q_{rc}}{1-2\cdot q_{rc}}$
$q_{cr}$ $=p\left[1-X|X,(y_{r},y_{c})\right]$
$q_{rc}$ $=p\left[1-X|X,(y_{c},y_{r})\right]$
where $q_{cr}$ and $q_{rc}$ are the probabilities that the LLM judge’s decision is flipped for both response orders.
In Appendix IV, we derive the above relationships, validate the position bias measurement based on de-noised accuracies, and provide a practical method for their computation.
Length Bias (LB) Previous studies have indicated that human evaluators exhibit the length bias when assessing responses [19, 22]. If LLM judges are employed as surrogates for human judges, it is expected they have the same length bias in general. Thus, this study aims to measure the relative length bias of LLM-judges compared with human evaluators, rather than their absolute length bias. For brevity, we use “length bias” to refer to the relative length bias in the paper.
For each data case $(x,\!y_{c},\!y_{r})$, we denote $\Delta l\!=\!l_{c}\!-\!l_{r}$ as the length difference between $y_{c}$ and $y_{r}$, where $l_{c}$ and $l_{r}$ are the length of $y_{c}$ and $y_{r}$, respectively. Additionally, we denote $p\left[X\!=\!1|\Delta l\!>0\right]$ as the probability that the LLM-judge’s result aligns with the human selection when the human selected response $y_{c}$ is longer than $y_{r}$, and $p\left[X\!=\!1|\Delta l\!\leq 0\right]$ as the probability that the LLM-judge’s result align when the length relationship is reversed. Moreover, these two probabilities are defined within the same accuracy framework, analogous to the definition of position bias.
Following the same rationale as in the position bias section, we define length bias as
$\text{LB}\!=p\left[X\!=\!1|\Delta l\!>0\right]\!-p\left[X\!=\!1|\Delta l\!\leq 0\right],$
[…]
In cases where flipping noise cannot be neglected, analogous to the approach for position bias, we first compute accuracies from noisy observations $Z$: $p\left[Z\!=\!1|\Delta l\!>0\right]$ and $p\left[Z\!=\!1|\Delta l\!\leq 0\right]$. We then apply a de-noising process based on the relationships between accuracy derived from $X$ and accuracy derived from $Z$ as follows:
$p\left[X=1|\Delta l>0\right]$ $=\frac{p\left[Z=1|\Delta l\!>0\right]-q_{\Delta l>0}}{1-2\cdot q_{\Delta l>0}}$
$p\left[X=1|\Delta l\leq 0\right]$ $=\frac{p\left[Z=1|\Delta l\!\leq 0\right]-q_{\Delta l\leq 0}}{1-2\cdot q_{\Delta l\leq 0}}$
$q_{\Delta l>0}$ $=p\left[1-X|X,\Delta l\!>0\right]$
$q_{\Delta l\leq 0}$ $=p\left[1-X|X,\Delta l\!\leq 0\right]$
where $q_{\Delta l>0}$ and $q_{\Delta l\leq 0}$ are the probabilities that the LLM judge’s decision is flipped for the conditions $\Delta l>0$ and $\Delta l\leq 0$, respectively.
In Appendix IV, we derive the above relationships, validate the length bias measurement based on de-noised accuracies, and provide a practical method for their computation.
Evaluation Framework
In this study, we introduce an evaluation framework that integrates our proposed methods for computing metrics, including accuracy ($\text{Acc}_{\text{both}}$, $\text{Acc}_{\text{random}}$), position bias and length bias. Furthermore, a set of visualization tools is developed to facilitate the analysis and comparison of the reliability of various LLM judges and their alignment with human preferences.
The pipeline of the framework, as depicted in Fig. 1, is structured into four modular components: 1) Data Sampler, 2) LLM Judges, 3) Metrics Computation, and 4) Metrics Visualization. The functionality of each component is detailed as follows.
Figure 1: LLM-as-a-Judge Evaluation Framework
Data Sampler: When dealing with a large human preference dataset and a limited budget for using commercial LLM models, it becomes necessary to sample a manageable-size subset from the full dataset for LLM judge evaluation. Our framework employs a stratified sampling strategy to ensure that the subset maintains the same proportion of different conditions (e.g. length difference distribution) as the original dataset.
LLM Judges: As defined in the Background section, an LLM judge refers to the combination of a particular LLM and a specific prompt template. Given an LLM judge, this module is responsible for generating textual judging decisions for each sampled data case and subsequently converting them into a binary outcome for metrics computation. This module allows the flexible creation of varied LLM judges by configuring different LLMs and prompt templates for evaluation.
Metrics Computation: This module computes alignment and reliability evaluation metrics (i.e. accuracy, position bias, and length bias) using the judging results from the LLM Judge module and the human preference labels provided by the dataset, based on the computational methods described in the Method section.
Metrics Visualization: This module visualizes both the individual computed metrics and their inter-relationships, providing comprehensive insights for comparing LLM judges and aiding in the selection of the most suitable LLM judge for specific LLM-alignment tasks.
Experiments
In this section, we introduce the datasets and LLM judges used in our experiments, as well as how to leverage them to compute the aforementioned evaluation metrics. The corresponding results and findings will be presented in the next section.
Data Selection
We demonstrate our evaluation framework using two datasets that are commonly used to evaluate LLM alignment algorithms: TL;DR summarization dataset [40, 41] and HH-RLHF-Helpfulness dataset [3]. Both datasets contain a prompt (a post for the summarization dataset and a conversation history between humans and LLM assistants for HH-RLHF dataset) with two responses generated by distinct LLMs for each sample. Also, human preference labels are available to indicate which response is more aligned with human preference. Both datasets have already been partitioned into train and test sets by the authors in the original studies.
In our experiment, it is highly time-consuming and expensive to evaluate LLM judges on all the data cases of both datasets (143,356 for summarization and 124,243 for HH-RLHF-Helpfulness), so we randomly sample a subset from each dataset to perform all the evaluation experiments. Compared with the summarization dataset, the HH-RLHF-Helpfulness dataset has a much smaller test set (6,240 vs. 70,228), thus, we select a subset from the TL;DR summarization test set following the previous study [5] and a subset from the entire HH-RLHF-Helpfulness dataset. Moreover, multiple data cases may share the same prompt (post or conversation history) with distinct response pairs. To make our collected datasets as diverse as possible, only one pair is kept for this prompt and others are removed. After this step, each unique prompt corresponds to only one unique answer pair. Then we randomly sample the prompts and their associated responses five times without replacement, resulting in five non-overlapping splits. Since measuring length bias requires dividing all the data cases into two conditions: whether longer responses are preferred by humans or not, we leverage the stratified sampling to preserve the same ratio of these two conditions as in the entire dataset.
Overall, both datasets used in our experiments contain 200 distinct samples for each split, which results in 1000 samples in total. The summarization and HH-RLHF-Helpfulness datasets have a stratified ratio (# of humans prefer longer responses: # of humans prefer shorter responses) of 115:85 and 111:89 respectively.
LLM Judges
Our LLM judges integrate a range of up-to-date and varied commercial large language models and prompt templates. Particularly, we assess GPT-4o, GPT-4o-mini and GPT-3.5-turbo with 8 templates on the summarization dataset and 10 templates on the HH-RLHF-Helpfulness dataset. Thus, there are $3\!\times\!8\!=\!24$ LLM judges for the summarization dataset and $3\!\times\!10\!=\!30$ LLM judges for the HH-RLHF-Helpfulness dataset.
GPT-4o is the most advanced model which has the latest checkpoint on 05/13/2024, GPT-4o-mini is the most cost-efficient model, while GPT-3.5-turbo is from the last OpenAI model generation and serves as the baseline in our experiments. Our preliminary studies suggest that GPT-4o exhibits comparable performance to GPT-4 in judging decision-making, but at a cost that is 4 to 6 times lower. Due to limited budget, we select GPT-4o for evaluation over GPT-4 from the list of commercial LLMs, despite GPT-4 being the most widely-used model in LLM alignment studies before the release of GPT-4o.
All the considered templates were actually used in the pairwise comparison tasks to evaluate different LLM alignment algorithms by papers of the last (2023) and this year (2024) and we make sure they all have dissimilar prompt formats. Furthermore, since our evaluation datasets have no “tied” labels from human annotations, which indicate two responses are equally preferred, we remove sentences from the prompt templates which allow LLM judges to select “tied” labels. Please refer to Table 1 and 2 for the complete list of all the prompt templates used in this study and the corresponding papers from which they are derived. Example templates for each dataset are provided in Appendix I.
Template Name Paper Link
Publication Time
guo Guo et al. [42] 02/2024
scheurer Scheurer et al. [43] 02/2024
liusie Liusie et al. [39] 02/2024
wang Wang et al. [13] 01/2024
zheng Zheng et al.[19] 12/2023
wu Wu et al. [44] 11/2023
chen Chen et al. [45] 09/2023
rafailov Rafailov et al. [5] 07/2023
Table 1: Prompt templates used for the TL;DR summarization dataset.
Template Name Paper Link
Publication Time
cheng Cheng et al. [45] 06/2024
zeng Zeng et al. [46] 04/2024
shen Shen et al. [47] 02/2024
guo Guo et al. [42] 02/2024
zheng Zheng et al.[19] 12/2023
mehta Mehta et al.[48] 12/2023
wu Wu et al. [44] 11/2023
bai Bai et al. [49] 11/2023
rafailov Rafailov et al. [5] 07/2023
xu Xu et al. [50] 05/2023
Table 2: Prompt templates used for the HH-RLHF-Helpfulness dataset.
Temperature Parameter Selection
Temperature parameter determines how deterministic LLM outputs are, which might affect the performance of LLM-judges. However, few previous studies that use LLMs as judges explicitly explain how and why they choose the temperature in their experiments. In this study, we assess the impact of the temperature parameter on the self-consistency (i.e. 1-flipping probability $q$) and accuracies of the large language models, which helps to select the temperature before evaluating LLM-judge performance using other metrics.
In detail, we investigate five temperature settings: $0.0$, $0.1$, $0.3$, $0.5$, and $0.7$. For each temperature setting, we concatenate data samples in all 5 splits (1000 samples in total) and repeatedly ask LLM judges to select the better response $K\!=\!5$ times for each sample. We compute the self-consistency for both response positions ($y_{c}$, $y_{r}$) and ($y_{r}$, $y_{c}$) separately, as well as $\text{Acc}_{\text{both}}$ across all the samples.
Through preliminary experiments, we found the impact of different temperatures is the same to the same LLM with different prompt templates, so in the large-scale experiments, only the prompt templates from DPO paper [5] are utilized for both datasets.
Metrics Computation
In this section, we introduce how to compute the flipping probability $q$ of flipping noise and other evaluation metrics in our experiments.
To compute the flipping probability, same as selecting the temperature parameter, we let LLM judges select their preferred response from each sample repeatedly for $K=5$ times. However, since we need to compute this probability for every LLM judge (24 for the summarization dataset and 30 for the HH-RLHF-Helpfulness dataset), we only leverage the first split of each dataset instead of all five splits due to limited budget. For each sample, the flipping probabilities $q_{cr}$ and $q_{rc}$ for both positions $(y_{c},y_{r})$ and $(y_{r},y_{c})$ are computed separately to estimate de-noised position bias, and the flipping probabilities $q_{\Delta l>0}$ and $q_{\Delta l\leq 0}$ are computed as well to calculate de-noised length bias.
To compute accuracy, position bias, and length bias, we compute each metric on all the splits ($S=5$). In the result, we report the mean and standard deviation of LLM judge performances across these five splits.
Results
In this section, we present the results and findings regarding temperature, accuracy, position bias, and length bias in our experiments.
Temperature
Table 3 contains the results of self-consistent rate (SCR) and accuracy with various temperatures. The self-consistent rate, given by $1-q$ as defined in Eq. 1, measures the probability that the LLM’s judgments are consistent across identical inputs. Since different LLMs show the same trend on both datasets, we only include GPT-4o here for the demonstration. Results regarding other LLMs are included in Appendix III.
From the table, we observe that higher temperatures result in lower self-consistency for both positions, while accuracy is not significantly affected by temperatures. Specifically, even when the temperature is set to 0.0, complete self-consistency (i.e. SCR=1.0) remains unachievable. Furthermore, self-consistency varies with different positions, thereby necessitating the separate measurement of flipping probabilities related to flipping noise associated with each position.
Finally, we aim to demonstrate the generalizability of our evaluation framework by employing a value that is not a special case, such as 0.0. Thus, we select 0.1 as the temperature in all of our experiments, which has the highest level of self-consistency compared with higher temperatures.
TL;DR Summarization HH-RLHF-Helpfulness
Temperature
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
0.0 0.977 0.971 0.665 (0.003) 0.974 0.967 0.573 (0.005)
0.1 0.973 0.967 0.666 (0.004) 0.966 0.957 0.575 (0.005)
0.3 0.963 0.956 0.668 (0.003) 0.950 0.944 0.574 (0.005)
0.5 0.953 0.949 0.663 (0.003) 0.942 0.926 0.579 (0.009)
0.7 0.946 0.927 0.657 (0.000) 0.934 0.914 0.577 (0.006)
Table 3: Self-consistent rate (SCR) and accuracy (Acc) of tested temperatures for the TL;DR summarization and HH-RLHF-Helpfulness datasets. Results are demonstrated using GPT-4o and prompt templates from the DPO paper [5].
Accuracy
Figure 2 shows accuracies ($\text{Acc}_{\text{both}}$) of LLM judges on both datasets, where identical colors represent the same prompt template within the same dataset (the same coloring rule is applied to all the result figures except for Figure 5). As we can see, different LLM judges have distinct accuracy, which means they have varied alignment levels with human preferences. Also, it demonstrates the performance of an LLM judge is highly sensitive to prompt templates.
Notably, several LLM judges have very low accuracies ($\text{Acc}_{\text{both}}\!<\!0.2$). Thus, it is significantly important to carefully evaluate and compare different LLM judges before actually using them to evaluate LLM alignment algorithms. Moreover, we find that all the accuracies on both datasets are below 0.7, which shows the mediocre alignment level and demonstrates that human evaluation is necessary to precisely compare different LLM alignment systems.
Compared with GPT-3.5-turbo, both GPT-4o and GPT-4o-mini have higher accuracies no matter which prompt template is used. It demonstrates that the superior internal capacities of recent LLMs, compared to older versions, are independent of the prompt templates used.
Figure 3 shows accuracy ($\text{Acc}_{\text{random}}$) of LLM judges on both datasets. Compared with Figure 2 (i.e. $\text{Acc}_{\text{both}}$), the gap between GPT-3.5-turbo and the others shrinks. This is because $\text{Acc}_{\text{random}}$ involves randomly selecting a position when LLM judge selection is inconsistent across two positions, thereby not reflecting the internal capabilities of LLM judges. Consequently, $\text{Acc}_{\text{random}}$ is a less effective metric for assessing LLM judge performance compared to $\text{Acc}_{\text{both}}$. Based on this, only $\text{Acc}_{\text{both}}$ is used to demonstrate the relationship between accuracy and position bias as well as length bias in the following sections.
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 2: Accuracy ($\text{Acc}_{\text{both}}$) for TL;DR the summarization and HH-RLHF-Helpfulness datasets. Please refer to Table 1 and Table 2 for details on the prompt templates used in all the result figures throughout the Results section. The results suggest the high sensitivity of LLM-judge accuracy to prompt templates and mediocre level of alignment to human judges.
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 3: Accuracy ($\text{Acc}_{\text{random}}$) for the TL;DR summarization and HH-RLHF-Helpfulness datasets. The comparison between results shown in Figure 2 and those shown in Figure 3 suggests that $\text{Acc}_{\text{random}}$ is a less effective metric for assessing LLM judge performance compared to $\text{Acc}_{\text{both}}$.
Position Bias
Position biases of all the LLM judges are shown in Figure 4, where positive values mean judges prefer the first position while negative values mean judges prefer the second position.
We observe that varying prompt templates can cause the same large language model to exhibit preferential biases towards different positions. Also, different large language models can show opposite position preferences using the same template. Thus, the position bias/preference depends on both the LLMs themselves and the prompt templates.
Additionally, we illustrate the relationship between accuracy and the absolute value of position bias in Figure 5. Here, absolute position bias reflects the bias level without specifying the preferred position. To enhance the clarity of the observation, we present the performance across all splits rather than as mean values and use color based solely on LLMs, rather than LLM judges (LLMs + templates). Our evaluation results reveal a significant negative correlation between accuracy and the level of position bias. The underlying reasons for this correlation need further investigation.
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 4: Position bias for the TL;DR summarization and HH-RLHF-Helpfulness datasets. Our results suggest that the position bias/preference depends on both the LLMs themselves and the prompt templates.
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 5: Absolute position bias vs. accuracy for TL;DR summarization and HH-RLHF-Helpfulness dataset. Our results reveal a significant negative correlation between accuracy and the level of position bias.
Length Bias
Figure 6 displays the (relative) length bias of all the judges across both datasets. Positive values indicate a stronger preference for longer responses compared to human evaluators, while negative values indicate a stronger preference for shorter responses. The figure shows that all the tested LLM judges have stronger preferences for longer responses compared to human judges, which is consistent with previous studies [19, 22]. Furthermore, compared to the summarization task, LLM judges exhibit a greater degree of length bias on the multi-turn conversation task (HH-RLHF-Helpfulness dataset).
Generally, longer responses tend to provide more detailed and comprehensive answers, which are more favored by humans compared to shorter ones [51, 52]. We suspect that the length bias results from the over-alignment of commercial models with human preferences.
Different from position bias, length bias does not have a negative correlation with accuracy (please refer to Appendix III for their relationship).
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 6: Length bias for TL;DR summarization and HH-RLHF-Helpfulness dataset. The results suggest that all the tested LLM judges have stronger preferences for longer responses compared to human judges, which might result from the over-alignment of the commercial models with human preferences.
Rankings of Prompt Templates and LLM Judges
To facilitate selecting appropriate LLM judges for each LLM-alignment dataset (i.e. TL;DR summarization and HH-RLHF-Helpfulness), we rank all the LLM judges (LLM + template) for each dataset, as well as all the prompt templates for each LLM used in our study (i.e. GPT-3.5-turbo, GPT-4o and GPT-4-mini) separately. We display top five templates or LLM judges and report their $\text{Acc}_{\text{both}}$, $\text{Acc}_{\text{random}}$, position bias and length bias. Please see Table 4 for the ranking results of all the LLM judges for the TL;DR summarization dataset and Table 5 for the HH-RLHF-Helpfulness dataset. The ranking results related to separate LLMs are provided in Appendix III.
Specifically, the rankings are based on $\text{Acc}_{\text{both}}$, which is because:
  • While position and length biases are critical metrics for assessing the reliability of LLM-based judges, accuracy is the metric that directly reflects their alignment with human preferences. Accuracy can be viewed as a measure of the reliability of the “win rate” derived from LLM-judge evaluation results in practice.
  • In the primary study, our findings indicate that $\text{Acc}_{\text{both}}$ more accurately represents the evaluative capabilities of LLM judges compared to $\text{Acc}_{\text{random}}$.
TL;DR Summarization (All LLMs)
Template / LLM $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
rafailov / gpt-4o 0.667 (0.011) 0.737 (0.014) 0.022 (0.015) 0.197 (0.031)
chen / gpt-4o 0.658 (0.028) 0.734 (0.029) -0.081 (0.023) 0.117 (0.055)
guo / gpt-4o 0.655 (0.011) 0.733 (0.024) -0.140 (0.014) 0.193 (0.038)
liusie / gpt-4o 0.632 (0.023) 0.724 (0.019) -0.154 (0.041) 0.084 (0.056)
rafailov / gpt-4o-mini 0.631 (0.014) 0.701 (0.023) -0.060 (0.027) 0.162 (0.038)
Table 4: Rankings of LLM judges (LLM+prompt template) on the TL;DR summarization dataset.
HH-RLHF-Helpfulness (All LLMs)
Template / LLM $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
guo / gpt-4o 0.618 (0.040) 0.694 (0.030) -0.005 (0.013) 0.135 (0.075)
xu / gpt-4o 0.610 (0.025) 0.702 (0.019) 0.086 (0.010) 0.029 (0.057)
bai / gpt-4o 0.603 (0.027) 0.697 (0.014) 0.034 (0.017) 0.255 (0.067)
guo / gpt-4o-mini 0.602 (0.036) 0.681 (0.030) -0.028 (0.026) 0.294 (0.059)
rafailov / gpt-4o-mini 0.594 (0.014) 0.657 (0.019) 0.047 (0.020) 0.463 (0.039)
Table 5: Rankings of LLM judges (LLM+prompt template) on HH-RLHF-Helpfulness dataset.
Limitations and Future Work
In this section, we discuss the limitations in this study and outline the directions for future research.
First, our current studies focus on commercial LLMs (e.g., GPT-3.5, GPT-4o, and GPT-4o-mini) rather than open-source LLMs. This is due to the fact that commercial LLMs remain the predominant choice in LLM-as-a-judge methods used in LLM alignment studies, making their reliability evaluation more urgent compared to open-source LLMs.
Second, our evaluation studies concentrate on LLM-as-a-judge methods, although open-source reward models (RMs) also hold the potential to serve as judges on LLM alignment tasks [53]. Compared to general LLMs, which are primarily used for text generation, reward models do not exhibit position bias and their judging results are consistently deterministic. Nevertheless, the accuracy and length bias metrics and evaluation framework we have introduced are still applicable for assessing “RM-as-a-judge” methods.
Lastly, while our evaluation results confirm the presence of position and length biases, which are commonly observed in LLM alignment studies, the accuracy of the metrics has not been thoroughly investigated. Comprehensive validation of the defined evaluation metrics would require extensive human-based assessments, which are not available in this study.
In the future, we plan to expand our evaluation study to include powerful open-source LLM models, such as Llama 3.1 [54], and open-source reward models, such as Nemotron-4-340B-Reward [53], across a broader range of datasets, including RewardBench [55]. Additionally, we will advance the evaluation of the introduced reliability metrics once human assessment resources become available.
Conclusions
In this study, we introduced a set of reliability metrics, including accuracy, position bias, and length bias, with improved theoretical interpretability. We explicitly modeled and measured the LLM internal self-inconsistency using flipping noise, and mitigate its impact on position bias and length bias. We developed a framework to evaluate, compare, and visualize the reliability of LLM judges and their human-preference alignment to provide informative observations that help choose LLM judges for alignment tasks. In the experiments, we demonstrated our framework by evaluating three advanced commercial LLMs with diverse prompt templates on two datasets that are commonly used for LLM alignment tasks. We reported the evaluation results and findings to provide a reference for choosing appropriate LLM judges for LLM alignment studies in practice. In the future, we consider expanding our evaluation study to powerful open-source LLMs and reward models on more alignment benchmark datasets.
Appendix
I. Prompt Templates in This Study
Examples of Prompt Templates
(TL;DR Summarization Dataset)
Template from Rafailov et al. [5]
Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details? A good summary is both precise and concise.
Post: <post>
Summary A: <summary A>
Summary B: <summary B>
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only "A" or "B" to indicate your choice. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <"A" or "B">
Template from Wang et al. [13]
As a neutral observer, your task is to assess the responses provided by two TL;DR summarizations according to the same SUBREDDIT prompt shown below. Begin by comparing the two responses and provide a brief explanation. Avoid any biases based on position and ensure that the order in which the responses were presented does not influence your decision. Do not let the length of the responses influence your evaluation. Do not favor certain names of the assistants. Strive to be as objective as possible. You need to choose only one of the two answers and respond by either A or B.
{prompt}
A. {answer_a}
B. {answer_b}
Which one is better? A or B?
Examples of Prompt Templates
(HH-RLHF-Helpfulness Dataset)
Template from Rafailov et al. [5]
For the following query to a chatbot, which response is more helpful?
Query: {the user query}
Response A:
{either the test method or baseline}
Response B:
{the other response}
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful. SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful: <"A" or "B">
Template from Shen et al. [47]
Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions better and provides more tailored responses to the user’s questions.
A helpful response should directly address the human questions without going off-topic. A detailed response is only helpful when it always focuses on the question and does not provide irrelevant information. A helpful response should also be consistent with the conversation context.
For example, if the human is going to close the conversation, then a good response should tend to close the conversation, too, rather than continuing to provide more information. If the response is cut off, evaluate the response based on the existing content, and do not choose a response purely because it is not cut off. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any positional biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor specific names of the assistants.
Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: [[A]] if assistant A is better, [[B]] if assistant B is better. Please make sure the last word is your choice.
--User Question--
{prompt}
--The Start of Assistant A’s Answer--
{response_1}
--The End of Assistant A’s Answer--
--The Start of Assistant B’s Answer--
{response_2}
--The End of Assistant B’s Answer--
II. Human Preference Data Used in This Study
Example from the TL;DR Summarization dataset
Post:
"SUBREDDIT: r/relationship_advice
TITLE: [17/m] in a sticky situation with her [17/f], my Asian parents, and the school administration
POST: Over two years ago my girlfriend and I started dating in secret. We were in secret because my parents are (racist?) in the way that they only want me to date people from an Asian background like me, and she is white. Eventually, because our school is super small and rumors spread like crazy, the staff found out maybe about a year ago. We went and made sure they knew not to go to our parents, and they all agreed. Fast forward to now and the principal and guidance counselor have called my parents and spilled the entire story to them. They apparently even had to use generic words like "girlfriend and her mom" instead of saying names to get around privacy rules. After talking it over with some of our close friends, no one has any insight or heard of anything that could cause them to do this, and it’s very uncharacteristic of them. My parents have told me that the school administration has said things such as, "She’s in a lower social class, he can do better," "She’s bringing his grades down" (I have a 4.0 GPA), etc. While my parents have also said things such as "She’s white trash," and "She’s a gold digger," I know for sure that those are both incorrect. Heck, my parents haven’t even spoken to her for more than 3 minutes. Now that the jig is up, I’m not sure what to do. Do I secretly talk to her and risk losing everything except food and school supplies for the next year if I get caught? Do I make arrangements to get back together in a year? Or Do I just give up and accept it as over? All of this stress is really getting to me, and I have no idea what the best choice is. Thanks for reading all of that, and giving me any feedback!
TL;DR: ’
Chosen Summary by Human Evaluators:
"[17/m] in a sticky situation with her [17/f], my Asian parents, and the school administration"
Rejected Summary by Human Evaluators:
"My parents found out my girlfriend and I are dating and the school administration has told them she is ’white trash’ and ’a gold digger’"
Example from the HH-RLHF-Helpfulness dataset
Single-turn Conversation Example
Conversation:
"Human: What are the rules to basketball?
Assistant: "
Chosen Response by Human Evaluators:
"Well, it’s played by two teams of five people each. Each team tries to get the ball through the other team’s hoop. The team that scores more points than their opponent is the winner."
Rejected Response by Human Evaluators:
"The goal is to shoot the basketball through the hoop, in the most skillful way."
Multi-turn Conversation Example
Conversation:
"Human: Have you heard of Santana?
Assistant: Yes! "Evil Ways", and "Smooth", right?
Human: I like "Smooth". Have you heard of it?
Assistant: Yes, it’s a great song! I like how the vocal is nice and slow at first, then faster at the end. It’s one of my favorite Santana songs.
Human: Someone else sang the song. Santana played guitar accompanying the song. Do you know the name of the singer?
Assistant: Oh yeah, I know who sang "Smooth". He’s a great singer. His name is Chris Standring.
Human: Both of your responses were inaccurate. The singer who sang "Smooth" is Rob Thomas.
Assistant: "
Chosen Response by Human Evaluators:
"Oh, I’m sorry. I didn’t recognize Rob Thomas. "Smooth" is a great song. Did you know Rob Thomas was a cast member of the band One Tree Hill?"
Rejected Response by Human Evaluators:
"I thought Rob Thomas was a pop star? I thought he played guitar and drums."
III. Additional Results
Temperature Results
TL;DR Summarization HH-RLHF-Helpfulness
Temperature
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
0.0 0.976 0.972 0.659 (0.002) 0.973 0.973 0.585 (0.002)
0.1 0.972 0.968 0.660 (0.003) 0.965 0.966 0.585 (0.003)
0.3 0.964 0.963 0.661 (0.006) 0.947 0.944 0.586 (0.003)
0.5 0.954 0.951 0.655 (0.003) 0.942 0.926 0.579 (0.009)
0.7 0.939 0.941 0.650 (0.004) 0.924 0.916 0.578 (0.008)
Table 6: Self-consistent rate (SCR) and accuracy (Acc) related to the tested temperatures for TL;DR summarization and HH-RLHF-Helpfulness datasets. Results are demonstrated using GPT-4o and the prompt template chen [56] for the summarization dataset and the template zeng [57] for the HH-RLHF-Helpfulness dataset, respectively. The conclusions are the same as those using prompt templates from the templates rafailov [5] for both datasets.
TL;DR Summarization HH-RLHF-Helpfulness
Temperature
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
0.0 0.989 0.991 0.631 (0.001) 0.987 0.990 0.589 (0.003)
0.1 0.986 0.985 0.630 (0.001) 0.983 0.988 0.591 (0.003)
0.3 0.974 0.982 0.627 (0.003) 0.970 0.968 0.593 (0.003)
0.5 0.972 0.978 0.629 (0.004) 0.965 0.967 0.587 (0.003)
0.7 0.961 0.973 0.622 (0.003) 0.960 0.957 0.585 (0.006)
Table 7: Self-consistent rate (SCR) and accuracy (Acc) related to the tested temperatures for TL;DR summarization and HH-RLHF-Helpfulness datasets. Results are demonstrated using GPT-4o-mini and prompt templates rafailov [5] for both datasets.
TL;DR Summarization HH-RLHF-Helpfulness
Temperature
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
SCR
($y_{c}$, $y_{r}$)
SCR
($y_{r}$, $y_{c}$)
Acc
($\text{Acc}_{\text{both}}$)
0.0 0.948 0.936 0.554 (0.004) 0.970 0.951 0.371 (0.003)
0.1 0.925 0.907 0.548 (0.008) 0.964 0.948 0.369 (0.002)
0.3 0.876 0.856 0.538 (0.003) 0.941 0.906 0.373 (0.006)
0.5 0.824 0.807 0.516 (0.008) 0.925 0.889 0.375 (0.010)
0.7 0.780 0.772 0.498 (0.006) 0.901 0.853 0.382 (0.008)
Table 8: Self-consistent rate (SCR) and accuracy (Acc) related to the tested temperatures for the TL;DR summarization and HH-RLHF-Helpfulness datasets. Results are demonstrated using GPT-3.5-turbo and prompt templates rafailov [5] for both datasets. GPT-3.5-turbo is much more sensitive to temperatures compared with GPT-4o and GPT-4o-mini.
Length Bias and Accuracy Relationship
Please refer to Figure 7 for the relationship between length bias and accuracy.
(a) TL;DR summarization
(b) HH-RLHF-Helpfulness
Figure 7: Length bias vs. accuracy for the TL;DR summarization and HH-RLHF-Helpfulness datasets.
Rankings of Prompt Templates and LLM Judges
Please see the ranking results of prompt templates for separate LLMs (i.e. GPT-3.5-turbo, GPT-4o, GPT-4o-mini) in Table 9 - 11 for TL;DR Summarization and Table 12 - 14 for HH-RLHF-Helpfulness dataset.
TL;DR Summarization (GPT-3.5-turbo)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
guo 0.566 (0.020) 0.675 (0.019) -0.047 (0.017) 0.174 (0.039)
rafailov 0.547 (0.022) 0.668 (0.040) 0.049 (0.018) 0.152 (0.045)
chen 0.516 (0.028) 0.652 (0.030) -0.291 (0.020) 0.085 (0.050)
liusie 0.496 (0.044) 0.654 (0.038) 0.204 (0.032) 0.237 (0.078)
wang 0.464 (0.023) 0.640 (0.027) 0.240 (0.039) 0.089 (0.094)
Table 9: Rankings of prompt templates for GPT-3.5-turbo on the TL;DR summarization dataset.
TL;DR Summarization (GPT-4o)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
rafailov 0.667 (0.011) 0.737 (0.014) 0.022 (0.015) 0.197 (0.031)
chen 0.658 (0.028) 0.734 (0.029) -0.081 (0.023) 0.117 (0.055)
guo 0.655 (0.011) 0.733 (0.024) -0.140 (0.014) 0.193 (0.038)
liusie 0.632 (0.023) 0.724 (0.019) -0.154 (0.041) 0.084 (0.056)
wang 0.601 (0.015) 0.695 (0.016) 0.108 (0.022) 0.137 (0.066)
Table 10: Rankings of prompt templates for GPT-4o on the TL;DR summarization dataset.
TL;DR Summarization (GPT-4o-mini)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
rafailov 0.631 (0.014) 0.701 (0.023) -0.060 (0.027) 0.162 (0.038)
guo 0.619 (0.032) 0.715 (0.010) 0.090 (0.036) 0.257 (0.068)
chen 0.615 (0.021) 0.692 (0.031) 0.010 (0.014) 0.104 (0.049)
liusie 0.563 (0.018) 0.684 (0.026) -0.122 (0.030) 0.169 (0.061)
zheng 0.516 (0.015) 0.667 (0.020) 0.280 (0.030) 0.544 (0.086)
Table 11: Rankings of prompt templates for GPT-4o-mini on the TL;DR summarization dataset.
HH-RLHF-Helpfulness (GPT-3.5-turbo)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
zeng 0.536 (0.012) 0.654 (0.023) 0.013 (0.036) 0.531 (0.044)
guo 0.506 (0.025) 0.651 (0.016) 0.029 (0.060) 0.280 (0.062)
bai 0.458 (0.022) 0.659 (0.032) 0.317 (0.033) 0.342 (0.043)
zheng 0.423 (0.018) 0.594 (0.009) 0.368 (0.035) 0.581 (0.051)
xu 0.386 (0.027) 0.622 (0.030) 0.488 (0.037) 0.309 (0.050)
Table 12: Rankings of prompt templates for GPT-3.5-turbo on the HH-RLHF-Helpfulness dataset.
HH-RLHF-Helpfulness (GPT-4o)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
guo 0.618 (0.040) 0.694 (0.030) -0.005 (0.013) 0.135 (0.075)
xu 0.610 (0.025) 0.702 (0.019) 0.086 (0.010) 0.029 (0.057)
bai 0.603 (0.027) 0.697 (0.014) 0.034 (0.017) 0.255 (0.067)
cheng 0.589 (0.029) 0.664 (0.029) 0.049 (0.020) 0.364 (0.082)
zeng 0.580 (0.034) 0.674 (0.027) 0.139 (0.023) 0.402 (0.090)
Table 13: Rankings of prompt templates for GPT-4o on the HH-RLHF-Helpfulness dataset.
HH-RLHF-Helpfulness (GPT-4o-mini)
Template $\text{Acc}_{\text{both}}$ $\text{Acc}_{\text{random}}$ Position Bias Length Bias
guo 0.602 (0.036) 0.681 (0.030) -0.028 (0.026) 0.294 (0.059)
rafailov 0.594 (0.014) 0.657 (0.019) 0.047 (0.020) 0.463 (0.039)
zeng 0.587 (0.031) 0.650 (0.029) 0.032 (0.022) 0.494 (0.061)
xu 0.580 (0.018) 0.681 (0.017) 0.036 (0.015) 0.272 (0.065)
bai 0.576 (0.033) 0.665 (0.022) -0.086 (0.010) 0.397 (0.061)
Table 14: Rankings of prompt templates for GPT-4o-mini on the HH-RLHF-Helpfulness dataset.
IV. Derivations, Proofs, and Computational Methods
Position Bias (PB)
1) Proof: Position bias definition is intrinsically length bias-mitigated
In this proof, we demonstrate that the impact of length bias has been effectively mitigated from the measurement of position bias using the definition in the main paper.
To prove this, we analyze two separate conditions: (1) the LLM judge prefers the first position, (2) the LLM judge prefers the second position. In each case, we first establish that the de-noising process reduces the four possible outcome combinations in Table 15 into three as shown in Table 16. Subsequently, we demonstrate that the measurement of position bias, utilizing de-noised accuracy, effectively mitigates the length bias.
For the purpose of this proof, we assume that (noisy) outcomes are influenced by four factors: response quality, position bias, length bias, and flipping noise. This assumption will be relaxed at the end of the proof. Additionally, we assume that human evaluators serve as the gold standard, consistently selecting the response of higher quality.
Before formally prove the claim, we remind readers that the position bias is defined based on the setting where the LLM judge decides on two reversed response orders for each data case: $h=(x,y_{c},y_{r})$ and $h^{\prime}=(x,y_{r},y_{c})$, which results in two outcomes $y$ and $y^{\prime}$ ($y,y^{\prime}\in\{y_{c},y_{r}\}$). Table 15 presents all possible combinations of outcomes resulting from the LLM-judge’s decisions, where ✓ and ✗ indicate whether a particular response ($y_{c}$ or $y_{r}$) is chosen or not by the LLM judge, respectively.
$y$ $y^{\prime}$
$y_{c}$ $y_{r}$ $y_{r}$ $y_{c}$
Table 15: All possible outcomes from LLM judge decisions.
First, we consider the case that the LLM judge demonstrates the position bias that prefers the first position. Consequently, we can examine the likely causes for each outcome $y,y^{\prime}=(y_{c},y_{r},y_{r},y_{c})$:
  • (✓✗✗✓): The LLM judge has selected the same response as human evaluators on both positions, either by emphasizing the response quality or due to the length bias (e.g. $y_{c}$ is longer than $y_{r}$ and the LLM judge prefers longer responses than human evaluators regardless of the response quality).
  • (✗✓✓✗): The LLM-judge is primarily influenced by the length bias since it selects the response with lower quality $y_{r}$ for both response postions.
  • (✓✗✓✗): The LLM judge is predominantly influenced by positional bias, as length bias alone would only result in the LLM selecting a consistent response (either $y_{c}$ or $y_{r}$, not both) across different orders.
  • (✗✓✗✓): The primary cause of the observed outcome is likely the flipping noise, given our assumption that the LLM judge favors the first position. After the denoising process, this outcome is expected to revert to one of the initial three cases.
  • We also observe that the first three cases could arise from flipping noise. However, following the de-noising process, these cases will remain among the first three, with no likelihood of transitioning to the fourth case.
Therefore, if the LLM judge exhibits the position bias towards the first position, the outcomes of the LLM-judge decisions with no flipping noise on $h$ and $h^{\prime}$ are shown in Table 16(a). Thus, the PB of the LLM judge is computed as:
$\text{PB}_{\text{first}}$ $\!=\!p\left[X=1|(y_{c},y_{r})\right]-p\left[X=1|(y_{r},y_{c})\right]$
$\!=\!\lim_{N\rightarrow\infty}\!\frac{1}{N}\sum_{n=1}^{N}\!\mathds{1}\!\left(y^{(n)}\!=\!y^{(n)}_{c}\right)\!-\!\lim_{N\rightarrow\infty}\!\frac{1}{N}\sum_{n=1}^{N}\!\mathds{1}\!\left(y^{\prime(n)}\!=\!y^{(n)}_{c}\right)$
[…]
$\!=\!\lim_{N\rightarrow\infty}\!\frac{1}{N}\sum_{n=1}^{N}\!\mathds{1}\left(y^{(n)}\!=\!y^{(n)}_{c}\land y^{\prime(n)}\!=\!y^{(n)}_{r}\right).$
[…]
Finally, if the observed outcomes are influenced by factors beyond response quality, positional bias, length bias, and flipping noise, these factors can be categorized into two types: position-dependent and position-independent. Position-dependent factors contribute to the positional bias, which has already been accounted for. Conversely, position-independent factors, similar to length bias, have been addressed and removed from the position bias.
Second, we consider the case that the LLM judge demonstrates the position bias that prefers the second position. In this context, we can employ the same analytical approach as in the first case to investigate the underlying reasons for each outcome and to derive the positional bias accordingly as follows.
$\text{PB}_{\text{second}}$ $\!=-\!\lim_{N\rightarrow\infty}\!\frac{1}{N}\sum_{n=1}^{N}\!\mathds{1}\left(y^{(n)}\!=\!y^{(n)}_{r}\land y^{\prime(n)}\!=\!y^{(n)}_{c}\right)$
[…]
$y$ $y^{\prime}$
$y_{c}$ $y_{r}$ $y_{r}$ $y_{c}$
(a) Prefer first position
$y$ $y^{\prime}$
$y_{c}$ $y_{r}$ $y_{r}$ $y_{c}$
(b) Prefer second position
Table 16: De-noised outcomes of the LLM judge’s decision in cases where the LLM judge favors the (a) first and (b) second responses, respectively. Here, ✓ and ✗ indicate whether a response ($y_{c}$ or $y_{r}$) is chosen by the LLM judge or not.
2) Derivations of de-noised position bias
The derivations related to the de-noising process of PB are provided as follows. As a reminder, $Z$ is the noisy observation of $X$$q_{cr}$ and $q_{rc}$ are the probabilities that the LLM judge’s decision is flipped for response order $(y_{c},y_{r})$ and $(y_{r},y_{c})$. Specifically,
$q_{cr}$ $=p\left[1-X|X,(y_{c},y_{r})\right],$
$q_{rc}$ $=p\left[1-X|X,(y_{r},y_{c})\right].$
In this study, we assume the flipping probability does not depend on the value of $X$, which needs further investigation. Based on this assumption, the relationship between the accuracy $p\left[X=1|(y_{c},y_{r})\right]$ and $p\left[Z=1|(y_{c},y_{r})\right]$ is derived as follows:
$p\left[Z=1|(y_{c},y_{r})\right]$ $=\overset{X~{}\text{is not flipped}}{\overbrace{p\left[X|X,(y_{c},y_{r})\right]\cdot p\left[X=1|(y_{c},y_{r})\right]}}$
$~{}~{}~{}~{}+\underset{X~{}\text{is flipped}}{\underbrace{p\left[1-X|X,(y_{c},y_{r})\right]\cdot p\left[X=0|(y_{c},y_{r})\right]}}$
$=(1\!-\!q_{cr})\!\cdot\!p\left[X\!=\!1|(y_{c},y_{r})\right]\!$
$~{}~{}~{}~{}~{}+\!q_{cr}\!\cdot\!(1\!-p\left[X\!=\!1|(y_{c},y_{r})\right])$
$=(1-2\cdot q_{cr})\cdot p\left[X=1|(y_{c},y_{r})\right]+q_{cr}$
Therefore,
$p\left[X=1|(y_{c},y_{r})\right]=\frac{p\left[Z=1|(y_{c},y_{r})\right]-q_{cr}}{1-2\cdot q_{cr}}$
Accordingly, the relationship between $p\left[X=1|(y_{r},y_{c})\right]$ and $p\left[Z=1|(y_{r},y_{c})\right]$ is:
$p\left[X=1|(y_{r},y_{c})\right]=\frac{p\left[Z=1|(y_{r},y_{c})\right]-q_{rc}}{1-2\cdot q_{rc}}$
3) Position bias computation procedure
Given a dataset $\mathcal{D}\!=\!\{h_{n}|n\!=\!1\!\ldots\!N\}$, a practical method for computing the PB related to an LLM judge is described as follows:
Step 1: Accuracy (based on $Z$) Computation
Since LLM judge evaluation results consistently contain flipping noise, even with the temperature parameter set to 0.0, we first calculate the accuracy for both response positions $(y_{c},y_{r})$ and $(y_{r},y_{c})$.
In order to achieve this, we employ the LLM judge to generate judging result on each data in $\mathcal{D}$ by considering two response orders: $h\!=\!(x,y_{c},y_{r})$ and $h^{\prime}\!=\!(x,y_{r},y_{c})$. The judge then selects the preferred response $y$ and $y^{\prime}$ from each order $h$ and $h^{\prime}$, where $y,y^{\prime}\!\in\!\{y_{c},y_{r}\}$.
Broadly, we denote the set of judging results as $\mathcal{J}\!=\!\{s_{n}|n=1\ldots N\}$, where each result $s_{n}\!=\!(y^{(n)},y^{\prime(n)})$ represents the selection outcome from both response orders, respectively, across all the data cases in the dataset $\mathcal{D}$. Then the accuracy for each position can be computed as follows:
$\hat{p}\left[Z\!=\!1|(y_{c},y_{r})\right]\!$ $=\!\frac{1}{N}\sum_{n=1}^{N}\mathds{1}\left(y^{(n)}\!=\!y_{c}^{(n)}\right),$
$\hat{p}\left[Z\!=\!1|(y_{r},y_{c})\right]\!$ $=\!\frac{1}{N}\sum_{n=1}^{N}\mathds{1}\left(y^{\prime(n)}\!=\!y_{c}^{(n)}\right).$
Step 2: Flipping Probability Estimation
Repeat the identical judging experiments in the Step 1 for extra $K\!-\!1$ times. These $K$ repetitions of identical judging experiments result in an extended judging result set $\mathcal{J}^{\prime}\!=\!\{s^{\prime}_{n}|n=1\ldots N\}$, where $s^{\prime}_{n}=\left(y^{(n)}_{1},y^{(n)}_{2},...,y^{(n)}_{K},y^{\prime(n)}_{1},y^{\prime(n)}_{2},...,y^{\prime(n)}_{K}\!\right)$. The flipping probabilities $q_{cr}$ and $q_{rc}$ for the position orders $(y_{c},y_{r})$ and $(y_{r},y_{c})$ are then computed by:
$\hat{q}_{cr}\!$ $=1\!-\!\frac{1}{N}\!\sum_{n=1}^{N}\!\left\{\!\frac{k^{(n)}_{cr}}{K}\cdot\frac{k^{(n)}_{cr}\!-\!1}{K\!-\!1}\!+\!\frac{K\!-\!k^{(n)}_{cr}}{K}\cdot\frac{\!K\!-\!k^{(n)}_{cr}\!-\!1}{K\!-\!1}\right\},$
$\hat{q}_{rc}\!$ $=1\!-\!\frac{1}{N}\!\sum_{n=1}^{N}\!\left\{\!\frac{k^{(n)}_{rc}}{K}\cdot\frac{k^{(n)}_{rc}\!-\!1}{K\!-\!1}\!+\!\frac{K\!-\!k^{(n)}_{rc}}{K}\cdot\frac{\!K\!-\!k^{(n)}_{rc}\!-\!1}{K\!-\!1}\right\}.$
where $k^{(n)}_{cr}\!=\!\sum^{K}_{k=1}\mathds{1}\!(y^{(n)}_{k}\!=\!y^{(n)}_{c})$ and $k^{(n)}_{rc}\!=\!\sum^{K}_{k=1}\mathds{1}\!(y^{\prime(n)}_{k}\!=\!y^{(n)}_{r})$ are the numbers of choosing the first response in $s^{\prime(n)}$ for the orders $(y_{c}^{(n)},y_{r}^{(n)}\!)$ and $(y_{r}^{(n)},y_{c}^{(n)})$, respectively.
Step 3: De-noising Process
The position bias is computed as follows:
$\text{PB}=\frac{\hat{p}\left[Z=1|(y_{c},y_{r})\right]-\hat{q}_{cr}}{1-2\cdot\hat{q}_{cr}}-\frac{\hat{p}\left[Z=1|(y_{r},y_{c})\right]-\hat{q}_{rc}}{1-2\cdot\hat{q}_{rc}}$
Length Bias (LB)
1) Proof: Length bias measurement is entangled with position bias
Here we demonstrate the entanglement between length bias (LB) and position bias (PB) in LB measurements.
Assume the LLM judge exhibits position bias, namely $\text{PB}=p\left[X\!=\!1|(y_{c},y_{r})\right]-p\left[X\!=\!1|(y_{r},y_{c})\right]\!\neq\!0$. Let $\text{LB}_{\text{cr}}$ and $\text{LB}_{\text{rc}}$ be length biases measured for response order $(y_{c},y_{r})$ and $(y_{r},y_{c})$ in all the data cases. Mathematically, they can be formulated as follows:
$\text{LB}_{cr}\!=$ $p\left[X\!=\!1|\Delta l\!>0,(y_{c},y_{r})\right]\!-p\left[X\!=\!1|\Delta l\!\leq 0,(y_{c},y_{r})\right],$
$\text{LB}_{rc}\!=$ $p\left[X\!=\!1|\Delta l\!>0,(y_{r},y_{c})\right]\!-p\left[X\!=\!1|\Delta l\!\leq 0,(y_{r},y_{c})\right].$
Due to the position bias, $p\left[X\!=\!1|\Delta l\!>0,(y_{c},y_{r})\right]\!\neq\!p\left[X\!=\!1|\Delta l\!>0,(y_{r},y_{c})\right]$ and $p\left[X\!=\!1|\Delta l\!\leq 0,(y_{c},y_{r})\right]\!\neq\!p\left[X\!=\!1|\Delta l\!\leq 0,(y_{r},y_{c})\right]$. Thus, generally $\text{LB}_{cr}\!\neq\!\text{LB}_{rc}$, and LB is dependent on the response order. The analysis above demonstrates that LB is generally entangled with PB in its measurement. In the next part, we introduce a method to approximate accuracies $p\left[X\!=\!1|\Delta l\!>0)\right]$ and $p\left[X\!=\!1|\Delta l\!\leq 0)\right]$ by mitigating the effect of PB.
2) Accuracy defintion selection
Previous work [19, 21] suggests both $\text{Acc}_{\text{both}}$ and $\text{Acc}_{\text{random}}$ (refer to the main paper for the definitions) can effectively mitigate the position bias in accuracy measurement. Here, we demonstrate that $\text{Acc}_{\text{both}}$ is the better choice than $\text{Acc}_{\text{random}}$ in terms of mitigating the influence of position bias for length bias measurement.
Without loss of generality, we assume the LLM judge has the position bias favoring the first response. The possible outcomes of $y^{\prime}$ and $y^{\prime}$ after the de-noising process can be thus found in Table 16(a).
[…]
When $\text{Acc}_{\text{random}}$ is used for accuracy, it depends on the proportion of both the first and the third case in Table 16(a) (the second case is not considered as it does not contribute to accuracy). This is because $\text{Acc}_{\text{random}}$ randomly selects $y$ and $y^{\prime}$ with a 50% probability, giving the third case a 50% chance of contributing to the correct selection for accuracy.
As previously discussed, the third case is primarily attributed to position bias and thus cannot fully mitigate the influence of positional bias, unlike $\text{Acc}_{\text{both}}$. Thus, in our study, $\text{Acc}_{\text{both}}$ is used to compute accuracy $p\left[X\!=\!1|\Delta l\!>0)\right]$ and $p\left[X\!=\!1|\Delta l\!\leq 0)\right]$ in our length bias computation procedures.
3) Length bias computation procedure
Given a dataset $\mathcal{D}\!=\!\{h_{n}|n\!=\!1\!\ldots\!N\}$, a practical method for computing LB related to an LLM judge is described as follows:
Step 1: Accuracy (based on $Z$) Estimation
First, we use the same way as for computing position bias to generate the judging result set $\mathcal{J}$. Then in order to compute the length bias, we divide the dataset $\mathcal{D}$ into two subsets of $\mathcal{D}$: $\mathcal{D}_{\Delta l>0}=\!\{h|\Delta l>0,h\in\mathcal{D}\}$, and $\mathcal{D}_{\Delta l\leq 0}=\!\{h|\Delta l\leq 0,h\in\mathcal{D}\}$ and also divide the judging result set $\mathcal{J}$ into two subsets of $\mathcal{J}$: $\mathcal{J}_{\Delta l>0}=\!\{s|\Delta l>0,s\in\mathcal{J}\}$ and $\mathcal{J}_{\Delta l\leq 0}=\!\{s|\Delta l\leq 0,s\in\mathcal{J}\}$. The accuracy based on $Z$ can then be computed as follows:
$\hat{p}\left[Z\!=\!1|\Delta l\!>\!0\right]\!$ $=\!\frac{1}{\left|\mathcal{J}_{\Delta_{l}>0}\right|}\!\sum_{s\in\mathcal{J}_{\Delta_{l}>0}}\!\mathds{1}\!\left(y\!=\!y_{c}\!\land\!y^{\prime}\!=\!y_{c}\right),$
$\hat{p}\left[Z\!=\!1|\Delta l\!\leq\!0\right]\!$ $=\!\frac{1}{\left|\mathcal{J}_{\Delta_{l}\leq 0}\right|}\!\sum_{s\in\mathcal{J}_{\Delta_{l}\leq 0}}\!\mathds{1}\!\left(y\!=\!y_{c}\!\land\!y^{\prime}\!=\!y_{c}\right).$
Step 2: Flipping Probability Estimation
Analogous to the PB computation procedure, we repeat the identical judging experiments for extra $K\!-\!1$ times to get the extended judging set $\mathcal{J}^{\prime}\!=\!\{s^{\prime}_{n}|n=1\ldots N\}$, where $s^{\prime}_{n}=\left(y^{(n)}_{1},y^{(n)}_{2},...,y^{(n)}_{K},y^{\prime(n)}_{1},y^{\prime(n)}_{2},...,y^{\prime(n)}_{K}\!\right)$. Subsequently, we divide $\mathcal{J}^{\prime}$ into two subsets: $\mathcal{J}^{\prime}_{\Delta l>0}=\!\{s^{\prime}|\Delta l>0,s^{\prime}\in\mathcal{J}^{\prime}\}$ and $\mathcal{J}^{\prime}_{\Delta l\leq 0}=\!\{s^{\prime}|\Delta l\leq 0,s^{\prime}\in\mathcal{J}^{\prime}\}$. The flipping probabilities $q_{\Delta l>0}$ and $q_{\Delta l\leq 0}$ is then computed as follows:
$\hat{q}_{\Delta>0}\!$ $=1\!-\!\frac{1}{N_{+}}\!\sum_{s^{\prime}\in\mathcal{J}^{\prime}_{\Delta_{l}>0}}\left\{\!\frac{k_{s^{\prime}}}{K}\!\cdot\!\frac{k_{s^{\prime}}\!-\!1}{K\!-\!1}\!+\!\frac{K\!-\!k_{s^{\prime}}}{K}\!\cdot\!\frac{\!K\!-\!k_{s^{\prime}}\!-\!1}{K\!-\!1}\right\},$
$\hat{q}_{\Delta\leq 0}\!$ $=1\!-\!\frac{1}{N_{-}}\!\sum_{s^{\prime}\in\mathcal{J}^{\prime}_{\Delta_{l}\leq 0}}\left\{\!\frac{k_{s^{\prime}}}{K}\!\cdot\!\frac{k_{s^{\prime}}\!-\!1}{K\!-\!1}\!+\!\frac{K\!-\!k_{s^{\prime}}}{K}\!\cdot\!\frac{\!K\!-\!k_{s^{\prime}}\!-\!1}{K\!-\!1}\right\},$
where $N_{+}\!=\!|\mathcal{J}^{\prime}_{\Delta_{l}>0}|$ and $N_{-}\!=\!\left|\mathcal{J}^{\prime}_{\Delta_{l}\leq 0}\right|$. Additionally, $k_{s^{\prime}}=\!\sum^{K}_{k=1}\mathds{1}\left(y_{k}\!=\!y_{c}\!\land\!y^{\prime}_{k}\!=\!y_{c}\right)$ represents the number of times that the LLM judge chooses $y_{c}$ in both position orders for any $s^{\prime}\in\mathcal{J}^{\prime}$, respectively.
Step 3: De-noising Process
The length bias is computed as follows:
$\text{LB}=\frac{\hat{p}\left[Z=1|\Delta l\!>0\right]-\hat{q}_{\Delta l>0}}{1-2\cdot\hat{q}_{\Delta l>0}}-\frac{\hat{p}\left[Z=1|\Delta l\!\leq 0\right]-\hat{q}_{\Delta l\leq 0}}{1-2\cdot\hat{q}_{\Delta l\leq 0}}.$